# Adaptively Bypassing Vision Transformer Blocks for Efficient Visual Tracking

### v1

@@ 6 Conclusion
In this work, we delve into the efficiency of ViT-based visual tracking methods, highlighting a previously disregarded insight: the importance of semantic features or relationships at a specific abstraction level depends on the target and the contextual nuances of its environment. We proposed a bypass decision module and a block sparsity loss to facilitate learning the adaptive bypassing of ViT blocks, thereby enhancing tracking efficiency. Additionally, we reduced the dimension of latent token representations by adapting a pruning technique to further enhance efficiency. Our method can substantially enhance tracking speed without compromising considerable precision, unveil new opportunities for visual tracking applications with limited computational resources. Notably, our method is straightforward and seamlessly integrable into existing ViT-based tracking frameworks. Extensive experiments confirm its effectiveness, demonstrating that our ABTrack achieves state-of-the-art performance across multiple benchmarks.

### v2

@@ 6 Conclusion
In this work, we delve into the efficiency of ViT-based visual tracking methods, highlighting a previously underappreciated insight: the importance of semantic features or relationships at a specific abstraction level depends on the target and the contextual nuances of its environment. We proposed a bypass decision module and a block sparsity loss to facilitate learning adaptively bypassing ViT blocks, thereby enhancing tracking efficiency. Additionally, we reduced the dimension of latent token representations by introducing a local ranking pruning strategy to further enhance efficiency. Our method can substantially enhance tracking speed without compromising considerable precision, unveil new opportunities for visual tracking applications with limited computational resources. Notably, our method is straightforward and seamlessly integrable into existing ViT-based tracking frameworks. Extensive experiments confirm its effectiveness, demonstrating that our ABTrack achieves state-of-the-art performance across multiple benchmarks.